OpenStudy (sasogeek):

In the United States, 43% of people wear a seat belt while driving. If two people are chosen at random, what is the probability that both of them wear a seat belt?

OpenStudy (anonymous):

probability that a person wear seat belt = 43% = 43/100 probability that both person wear seat belt = 43/100 * 43/100 =0. 184 9 =18.49 %